# Why Rectified Power Unit Networks Fail and How to Improve It: An Effective Theory Perspective

### v3

@@ 3.1 Susceptibility Calculation
Activation functions that exhibit similar behaviors according to the RG flow are classified into universality classes. Universality class is a concept in statistical physics used to classify systems that exhibit similar behavior at the limit scale. According to [18], scale-invariant activations (e.g. ReLU, LeakyReLU) form a single universality class called the self-invariant universality class. Tanh and Sin fall into the $K^{\star}=0$ universality class, while SWISH and GELU belong to the half-stable universality class. On the other hand, Perceptron, Sigmoid, and Softplus are known to not satisfy criticality at all. We will now calculate the susceptibility for the RePU activation function following equation (5) and demonstrate that RePU also fails to satisfy criticality. For the RePU activation function of order $p$, the parallel susceptibility and perpendicular susceptibility are calculated as follows:
[…]
The parallel susceptibility to perpendicular susceptibility ratio is $2p-1:p$. Therefore, except for $p=1$ (i.e., ReLU), it is impossible to set both susceptibilities to 1 simultaneously. If we set the parallel susceptibility to 1, the perpendicular susceptibility will be less than 1, causing the distance between two inputs to vanish as they propagate through the layers. Conversely, if we set the perpendicular susceptibility to 1, the parallel susceptibility will be greater than 1, causing each input to diverge as it propagates through the layers.
@@ 3.2 Experimental Validation
Empirical Kernel Behavior at Initialization According to the results in section 3.1, the kernel value of Neural Networks with RePU (for $p>1$) for a single input must either explode or vanish as it propagates through the layers under the initialization distribution. In this section, we experimentally verify this. For the experiment, the neural network is configured with an input dimension of 2, an output dimension of 1, hidden layers with widths of 512, and a depth of 10 (therefore, the total number of layers, including the input and output layers, is 12). The activation function is RePU. The initialization distribution follows equation (2), with $C_{b}$ set to 0, and $C_{W}$ adjusted during the experiment. To experimentally observe the change in the kernel, we define the following empirical kernel quantities:
@@
Where $z^{(l)}_{i}(\Theta)|_{\mathcal{D}=x_{0}}$ is a random variable conditioned on the input $x_{0}$ and has randomness on $\Theta$. Similarly, $z^{(l)}_{i}(\mathcal{D})|_{\Theta=\Theta_{0}}$ is a random variable conditioned on the parameter $\Theta$ and has randomness on the data input. Each component of sample in $\mathcal{D}$ follows $N(0,1)$. $x_{0}$ is also generated from $\mathcal{D}$. The experimental results for a width of 512, $p=2$, and $C_{W}=21.3$ are shown in Figures 1. We observed that if $C_{W}$ is set smaller, the kernel vanishes on an $\mathcal{O}(\exp(\exp(-l)))$ scale, and if $C_{W}$ is set larger, it explodes on an $\mathcal{O}(\exp(\exp(l)))$ scale. The results in the figure for $C_{W}=21.3$ show that the kernel is highly sensitive to initial conditions and tends to either double exponentially explode or vanish. According to equations (5), (6) the susceptibility itself acts as a geometric ratio and is proportional to the kernel. As this ratio increases exponentially, it causes the kernel to grow double exponentially.
@@
Through this, we can infer that in RePU neural networks, the kernel either increases or decreases double exponentially, and since the perpendicular susceptibility is proportional to this, the training dynamics become highly unstable or converge rapidly. To experimentally verify this, we designed an experiment using synthetic data. The task involved observing how well the neural network performs regression on synthetic data, where the input data are points in the plane $(x,y)$ with $x\sim N(0,1)$ and $y\sim N(0,1)$, and the target is set to $x^{2}+y^{2}$. The training dataset is composed of 100 samples. We conducted experiments with a network width of 512 and with 1, 3, and 5 hidden layers, testing cases where $p=2$ and $p=3$. The $C_{W}$ values were adjusted based on the architecture; if $C_{W}$ is too small, the kernel decreases double exponentially, and if $C_{W}$ is too large, the kernel becomes highly unstable and eventually diverges. Therefore, we experimentally determined and set an appropriate $C_{W}$ value. We calculated the kernel for an ensemble of 100 neural networks, obtaining the mean and variance while excluding models where the kernel overflowed due to explosion. The results are presented in Tables 2 and 3, with the results for 5 hidden layers and $p=2$ depicted in Figure 2. We then tested the ensemble of 100 neural network models trained under each setting on a test dataset generated from the same distribution as the training data. The results are depicted in Figure 3. From the aggregated results of the kernel and the mean outputs on the test dataset, we observed that significant learning only occurs when there is a single hidden layer. When the network has 3 or 5 hidden layers, the kernel values quickly diminish as the layers deepen, leading to poor learning performance. Additionally, as the depth of the network increases, the mean output converges to a constant, with its standard deviation also converging to a very small value.
Epochs $(N_{h},C_{W})$ $(1,2.5)$ $(3,3.0)$ $(5,2.6)$ At Initialization $(1.39\text{*}10^{-3},3\text{*}10^{-4})$ $(1.28\text{*}10^{-6},1.5\text{*}10^{-6})$ $(6.91\text{*}10^{-11},1.9\text{*}10^{-10})$ At 250 Epoch $(1.71\text{*}10^{-3},4\text{*}10^{-4})$ $(2.80\text{*}10^{-5},1.2\text{*}10^{-5})$ $(4.29\text{*}10^{-7},2,7\text{*}10^{-7})$ At 500 Epoch $(2.71\text{*}10^{-3},8\text{*}10^{-4})$ $(7.44\text{*}10^{-5},5.3\text{*}10^{-5})$ $(1.03\text{*}10^{-6},7.2\text{*}10^{-7})$ At 750 Epoch $(4.57\text{*}10^{-3},1.5\text{*}10^{-3})$ $(1.35\text{*}10^{-4},9.7\text{*}10^{-5})$ $(1.75\text{*}10^{-6},1.3\text{*}10^{-6})$ At 1000 Epoch $(6.70\text{*}10^{-3},1.6\text{*}10^{-3})$ $(2.14\text{*}10^{-4},1.4\text{*}10^{-4})$ $(2.62\text{*}10^{-6},2.1\text{*}10^{-6})$
@@
Epochs $(N_{h},C_{W})$ $(1,4.5)$ $(3,2.9)$ $(5,2.7)$ At Initialization $(9.01\text{*}10^{-3},4\text{*}10^{-3})$ $(3.49\text{*}10^{-10},1.3\text{*}10^{-9})$ $(1.22\text{*}10^{-14},8.3\text{*}10^{-14})$ At 250 Epoch $(1.05\text{*}10^{-2},4\text{*}10^{-3})$ $(1.34\text{*}10^{-7},1.6\text{*}10^{-7})$ $(1.54\text{*}10^{-12},9.0\text{*}10^{-12})$ At 500 Epoch $(1.14\text{*}10^{-2},5.1\text{*}10^{-3})$ $(2.30\text{*}10^{-7},2.2\text{*}10^{-7})$ $(1.91\text{*}10^{-12},1.0\text{*}10^{-11})$ At 750 Epoch $(1.24\text{*}10^{-2},6.1\text{*}10^{-3})$ $(3.31\text{*}10^{-7},2.9\text{*}10^{-7})$ $(2.35\text{*}10^{-12},1.2\text{*}10^{-11})$ At 1000 Epoch $(1.36\text{*}10^{-2},7.5\text{*}10^{-3})$ $(4.53\text{*}10^{-7},4.0\text{*}10^{-7})$ $(2.81\text{*}10^{-12},1.5\text{*}10^{-11})$
@@ 4 Modified Rectified Power Unit (MRePU)
Definition 6 (Modified Rectified Power Unit (MRePU) The MRePU activation function of order $p>1$ is defined by the following equation:
@@
Similar to RePU, the MRePU function of order $p$ has the property of being differentiable $p-1$ times.
4.1 Numerical Calculation of Susceptibility
Here, we calculate the parallel and perpendicular susceptibilities of the MRePU function we proposed. The susceptibility of the MRePU function of order p can be expressed by the following equation:
[…]
Since getting value of (7) analytically is impossible, we calculate the susceptibility numerically. We fixed the value of $C_{W}$ at $1.0$ and plotted graphs by exponentially increasing the value of $K$ for cases where $p=2$ and $p=3$. The results can be seen in Figure 4, where it is observed that the ratio of the two susceptibilities converges to 1 as the value of $K$ approaches 0. This confirms that the MRePU activation belongs to the $K^{\star}=0$ universality class, and by appropriately adjusting the $C_{W}$ value, it is possible to ensure that the kernel neither explodes nor vanishes.
Figure 4: Parallel and Perpendicular Susceptibility and Their Ratio for MRePU with $p=2$ (Left) and $p=3$ (Right) with $(C_{W},C_{b})=(1,0)$
@@ 4.2 Experimental Results
Empirical Kernel and Four-point Vertex Behavior at Initialization As with the RePU case, we investigate the statistical properties of the neural network ensemble given by the initialization distribution. The experimental environment remains identical to that of the RePU Networks, with the only change being the activation function, which is now MRePU with $p=2$. The results for $C_{W}=1.0$ are shown in Figure 5. As can be seen from the graph, the empirical kernel maintains a consistent scale across both random data and weight parameters. Additionally, we also examined the behavior of the Four-point vertex for MRePU. Equation (4) can be simplified for a single input as follows:
@@
Figure 6 illustrates how the empirical four-point vertex, calculated from an ensemble of 1000 models using equation (3), changes across layers (with input $x_{0}=(1,0)$). $\mathcal{X}_{\parallel}$ increases as the kernel grows, and since the kernel itself depends on $C_{W}$, the inductive change in the four-point vertex is determined by $C_{W}$ according to equation (8). As evidence of this, the left figure shows that the four-point vertex increases more steeply as $C_{W}$ becomes larger. We also investigated the behavior of the four-point vertex across different widths. The results indicate that with smaller widths, perturbative effects in equation (8) become more pronounced, leading to a more rapid increase in the four-point vertex, and overall, higher values are observed. The four-point vertex serves as an indicator of non-Gaussianity, which is expected to increase as more layers with non-linear activations are added. Moreover, since smaller widths imply stronger interactions between neurons, these experimental results align well with our theoretical predictions.
@@
Dynamics of Statistics During Training
Now, we will examine how the statistics of the MRePU Network ensemble change under training dynamics. The experimental setup is the same as in the case of the RePU Network, and the number of hidden layers is set to 5. In this case, we constructed the training dataset in the same way as in the RePU activation case, but the number of samples is set to 1000. Although the theoretical approximation suggests setting $C_{W}=1.0$, meaningful learning was only achieved with a slightly lower value, possibly due to the perturbative effects in the MRePU Network. In our experiments, we set $C_{W}=0.8$. The dynamics of the kernel and four-point vertex are depicted in Figure 7, while the comparison of the ensemble’s mean output against the target for the test dataset is shown in Figure 8. Surprisingly, we observed that even with a vanilla FCN (Fully Connected Network) without considering architectures like ResNet, which are designed to prevent issues like the vanishing gradient, the deep neural network with 5 hidden layers still achieved meaningful learning using the MRePU function we designed. Additionally, as shown on the left side of Figure  7, we also calculated the perpendicular susceptibility, which showed a tendency to approach 1 as training progressed. On the right side of the same figure, it can be seen that the ensemble’s four-point vertex decreases to 0 during initialization but tends to increase linearly with the layers as training progresses. An increasing four-point vertex indicates growing interactions between neurons, which implies that representation learning is occurring. This suggests that the deep vanilla MRePU network is indeed engaging in meaningful representation learning.

### v4

@@ 3.1 Susceptibility Calculation
Activation functions that exhibit similar behaviors according to the RG flow are classified into universality classes. Universality class is a concept in statistical physics used to classify systems that exhibit similar behavior at the limit scale. According to (Roberts22), scale-invariant activations (e.g. ReLU, LeakyReLU) form a single universality class called the self-invariant universality class. Tanh and Sin fall into the $K^{\star}=0$ universality class, while SWISH and GELU belong to the half-stable universality class. On the other hand, Perceptron, Sigmoid, and Softplus are known to not satisfy criticality at all. We will now calculate the susceptibility for the RePU activation function following Eq. (6) and demonstrate that RePU also fails to satisfy criticality. For the RePU activation function of order $p$, the parallel susceptibility and perpendicular susceptibility are calculated as follows:
$\begin{split}\mathcal{X}_{\parallel}(K)&=\frac{C_{W}}{K}\langle z\sigma^{\prime}(z)\sigma(z)\rangle_{K}=\frac{C_{W}}{K}\langle z\lfloor pz^{p-1}\rfloor_{+}\lfloor z^{p}\rfloor_{+}\rangle_{K}\\ &=\frac{C_{W}p}{K\sqrt{2\pi K}}\int_{0}^{\infty}e^{-\frac{z^{2}}{2K}}z^{2p}dz=\frac{C_{W}p(2p-1)!!K^{p-1}}{2}.\end{split}$
$\begin{split}\mathcal{X}_{\perp}(K)&=C_{W}\langle\sigma^{\prime}(z)\sigma^{\prime}(z)\rangle_{K}=C_{W}\langle p^{2}\lfloor z^{p-1}\rfloor_{+}\lfloor z^{p-1}\rfloor_{+}\rangle_{K}\\ &=\frac{C_{W}p^{2}}{\sqrt{2\pi K}}\int_{0}^{\infty}e^{-\frac{z^{2}}{2K}}z^{2p-2}dz=\frac{C_{W}p^{2}(2p-3)!!K^{p-1}}{2}.\end{split}$ (7)
Here the notation $n!!$ means double factorial (e.g. $7!!=7\cdot 5\cdot 3\cdot 1$). The parallel susceptibility to perpendicular susceptibility ratio is 2p1:p2p-1:p. Therefore, except for $p=1$ (i.e., ReLU), it is impossible to set both susceptibilities to 1 simultaneously. If we set the parallel susceptibility to 1, the perpendicular susceptibility will be less than 1, causing the distance between two inputs to vanish as they propagate through the layers. Conversely, if we set the perpendicular susceptibility to 1, the parallel susceptibility will be greater than 1, causing each input to diverge as it propagates through the layers.
@@ 3.2 Experimental Validation
Empirical Kernel Behavior at Initialization According to the results in section 3.1, the kernel value of Neural Networks with RePU (for $p>1$) for a single input must either explode or vanish as it propagates through the layers under the initialization distribution. In this section, we experimentally verify this. For the experiment, the neural network is configured with an input dimension of 2, an output dimension of 1, hidden layers with widths of 512, and a depth of 10 (therefore, the total number of layers, including the input and output layers, is 12). The activation function is RePU. The initialization distribution follows Eq. (2), with $C_{b}$ set to 0, and $C_{W}$ adjusted during the experiment. To experimentally observe the change in the kernel, we define the following empirical kernel quantities:
@@
Where $z^{(l)}_{i}(\Theta)|_{\mathcal{D}=x_{0}}$ is a random variable conditioned on the input $x_{0}$ and has randomness on $\Theta$. Similarly, $z^{(l)}_{i}(\mathcal{D})|_{\Theta=\Theta_{0}}$ is a random variable conditioned on the parameter $\Theta$ and has randomness on the data input. Each component of sample in $\mathcal{D}$ follows $N(0,1)$. $x_{0}$ is also generated from $\mathcal{D}$. The experimental results for a width of 512, $p=2$, and $C_{W}=21.3$ are shown in Figs. 1. We observed that if $C_{W}$ is set smaller, the kernel vanishes on an $\mathcal{O}(\exp(\exp(-l)))$ scale, and if $C_{W}$ is set larger, it explodes on an $\mathcal{O}(\exp(\exp(l)))$ scale. The results in the figure for $C_{W}=21.3$ show that the kernel is highly sensitive to initial conditions and tends to either double exponentially explode or vanish. According to Eqns.(6) and (7) the susceptibility itself acts as a geometric ratio and is proportional to the kernel. As this ratio increases exponentially, it causes the kernel to grow double exponentially.
@@
Through this, we can infer that in RePU neural networks, the kernel either increases or decreases double exponentially, and since the perpendicular susceptibility is proportional to this, the training dynamics become highly unstable or converge rapidly. To experimentally verify this, we designed an experiment using synthetic data. The task involved observing how well the neural network performs regression on synthetic data, where the input data are points in the plane $(x,y)$ with $x\sim N(0,1)$ and $y\sim N(0,1)$, and the target is set to $x^{2}+y^{2}$. The training dataset is composed of 100 samples. We conducted experiments with a network width of 512 and with 1, 3, and 5 hidden layers, testing cases where $p=2$ and $p=3$. The $C_{W}$ values were adjusted based on the architecture; if $C_{W}$ is too small, the kernel decreases double exponentially, and if $C_{W}$ is too large, the kernel becomes highly unstable and eventually diverges. Therefore, we experimentally determined and set an appropriate $C_{W}$ value. We calculated the kernel for an ensemble of 100 neural networks, obtaining the mean and variance while excluding models where the kernel overflowed due to explosion. The results are presented in Tables 2 and 3, with the results for 5 hidden layers and $p=2$ depicted in Fig. 2. We then tested the ensemble of 100 neural network models trained under each setting on a test dataset generated from the same distribution as the training data. The results are depicted in Fig. 3. From the aggregated results of the kernel and the mean outputs on the test dataset, we observed that significant learning only occurs when there is a single hidden layer. When the network has 3 or 5 hidden layers, the kernel values quickly diminish as the layers deepen, leading to poor learning performance. Additionally, as the depth of the network increases, the mean output converges to a constant, with its standard deviation also converging to a very small value.
Epochs $(N_{h},C_{W})$ $(1,2.5)$ $(3,3.0)$ $(5,2.6)$ At initialization $(1.39\times 10^{-3},\,3\times 10^{-4})$ $(1.28\times 10^{-6},\,1.5\times 10^{-6})$ $(6.91\times 10^{-11},\,1.9\times 10^{-10})$ At 250 epochs $(1.71\times 10^{-3},\,4\times 10^{-4})$ $(2.80\times 10^{-5},\,1.2\times 10^{-5})$ $(4.29\times 10^{-7},\,2.7\times 10^{-7})$ At 500 epochs $(2.71\times 10^{-3},\,8\times 10^{-4})$ $(7.44\times 10^{-5},\,5.3\times 10^{-5})$ $(1.03\times 10^{-6},\,7.2\times 10^{-7})$ At 750 epochs $(4.57\times 10^{-3},\,1.5\times 10^{-3})$ $(1.35\times 10^{-4},\,9.7\times 10^{-5})$ $(1.75\times 10^{-6},\,1.3\times 10^{-6})$ At 1000 epochs $(6.70\times 10^{-3},\,1.6\times 10^{-3})$ $(2.14\times 10^{-4},\,1.4\times 10^{-4})$ $(2.62\times 10^{-6},\,2.1\times 10^{-6})$
@@
[…]
@@ 4 Modified Rectified Power Unit (MRePU)
Definition 7 (Modified Rectified Power Unit (MRePU). The MRePU activation function of order $p>1$ is defined by the following equation:
@@
Similar to RePU, the MRePU function of order $p$ has the property of being differentiable p1p-1 times.
4.1 Susceptibility Calculation
We compute the parallel and perpendicular susceptibilities of the proposed MRePU activation. For an MRePU of order $p$, they are given by
$\begin{split}\mathcal{X}_{\parallel}(K)&=\frac{C_{W}}{K}\,\big\langle z\,\sigma^{\prime}(z)\sigma(z)\big\rangle_{K}=\frac{C_{W}}{K\sqrt{2\pi K}}\int_{-1}^{\infty}(z+1)^{2p-1}\big((p+1)z+1\big)z^{2}\,e^{-z^{2}/(2K)}\,\mathrm{d}z,\\[5.69054pt] \mathcal{X}_{\perp}(K)&=C_{W}\,\big\langle\sigma^{\prime}(z)\sigma^{\prime}(z)\big\rangle_{K}=\frac{C_{W}}{\sqrt{2\pi K}}\int_{-1}^{\infty}(z+1)^{2p-2}\big((p+1)z+1\big)^{2}\,e^{-z^{2}/(2K)}\,\mathrm{d}z.\end{split}$ (8)
Proposition 8. Let $\phi(t)=\frac{1}{\sqrt{2\pi}}e^{-t^{2}/2}$ and $I_{n}(\alpha):=\int_{\alpha}^{\infty}y^{n}\phi(y)\,\mathrm{d}y$ with $\alpha=-1/\sqrt{K}$. Then
$\begin{split}\mathcal{X}_{\parallel}(K)&=C_{W}\sum_{n=2}^{2p+2}A_{n}(p)\,K^{\frac{n}{2}-1}\,I_{n}\!\Big(-\frac{1}{\sqrt{K}}\Big),\\ \mathcal{X}_{\perp}(K)&=C_{W}\sum_{n=0}^{2p}B_{n}(p)\,K^{\frac{n}{2}}\,I_{n}\!\Big(-\frac{1}{\sqrt{K}}\Big),\end{split}$ (9)
where, with the convention $\binom{m}{r}=0$ if $r\notin\{0,\dots,m\}$,
$A_{n}(p)=(p+1)\binom{2p-1}{\,n-3\,}+\binom{2p-1}{\,n-2\,},\qquad B_{n}(p)=(p+1)^{2}\binom{2p-2}{\,n-2\,}+2(p+1)\binom{2p-2}{\,n-1\,}+\binom{2p-2}{\,n\,}$.
Proof. (1) Polynomial expansion. Write
$(z+1)^{2p-1}\big((p+1)z+1\big)z^{2}=(p+1)\sum_{j=0}^{2p-1}\binom{2p-1}{j}z^{j+3}+\sum_{j=0}^{2p-1}\binom{2p-1}{j}z^{j+2}.$
Collecting the $z^{n}$ terms gives the coefficient in front of $z^{n}$ as $A_{n}(p)=(p+1)\binom{2p-1}{n-3}+\binom{2p-1}{n-2}$. Similarly,
$(z+1)^{2p-2}\big((p+1)z+1\big)^{2}=(p+1)^{2}\sum_{j=0}^{2p-2}\binom{2p-2}{j}z^{j+2}+2(p+1)\sum_{j=0}^{2p-2}\binom{2p-2}{j}z^{j+1}+\sum_{j=0}^{2p-2}\binom{2p-2}{j}z^{j},$
whence the $z^{n}$ coefficient is $B_{n}(p)=(p+1)^{2}\binom{2p-2}{n-2}+2(p+1)\binom{2p-2}{n-1}+\binom{2p-2}{n}$.
(2) Reduction to $I_{n}$. For any integer $n\geq 0$, the change of variables $y=z/\sqrt{K}$ yields
$\int_{-1}^{\infty}z^{n}\frac{e^{-z^{2}/(2K)}}{\sqrt{2\pi K}}\,\mathrm{d}z=K^{n/2}\int_{-1/\sqrt{K}}^{\infty}y^{n}\phi(y)\,\mathrm{d}y=K^{n/2}I_{n}\!\Big(-\frac{1}{\sqrt{K}}\Big).$
Combining (1)–(2) and summing termwise gives the stated formulas. $\square$
Corollary 9 (small-$K$ limit). As $K\to 0$,
$\mathcal{X}_{\parallel}(K)\;\longrightarrow\;C_{W},\qquad\mathcal{X}_{\perp}(K)\;\longrightarrow\;C_{W}.$
Proof. Let $\beta=1/\sqrt{K}\to\infty$. Since $I_{n}(-\beta)=\int_{-\beta}^{\infty}y^{n}\phi(y)\,\mathrm{d}y$ and $\phi$ is integrable with polynomial weights, dominated convergence gives
$\lim_{\beta\to\infty}I_{n}(-\beta)=\int_{-\infty}^{\infty}y^{n}\phi(y)\,\mathrm{d}y=\begin{cases}0,&n\ \text{odd},\\ (n-1)!!,&n\ \text{even}.\end{cases}$
In $\mathcal{X}_{\parallel}(K)=C_{W}\sum_{n=2}^{2p+2}A_{n}(p)\,K^{\frac{n}{2}-1}I_{n}(-\beta)$, all terms with $n>2$ vanish because $K^{\frac{n}{2}-1}\to 0$, while for $n=2$ we have $K^{0}=1$, $I_{2}(-\beta)\to 1$, and $A_{2}(p)=(p+1)\binom{2p-1}{-1}+\binom{2p-1}{0}=1$. Hence $\lim_{K\to 0}\mathcal{X}_{\parallel}(K)=C_{W}$. Likewise, in $\mathcal{X}_{\perp}(K)=C_{W}\sum_{n=0}^{2p}B_{n}(p)\,K^{\frac{n}{2}}I_{n}(-\beta)$, the leading term is $n=0$ with $K^{0}=1$, $I_{0}(-\beta)\to 1$, and $B_{0}(p)=(p+1)^{2}\binom{2p-2}{-2}+2(p+1)\binom{2p-2}{-1}+\binom{2p-2}{0}=1$; all $n\geq 1$ terms vanish. Thus $\lim_{K\to 0}\mathcal{X}_{\perp}(K)=C_{W}$. $\square$
Numerics.
For visualization we evaluate the above closed forms numerically. We fix $C_{W}=1$ and sweep $K$ on a logarithmic grid for $p\in\{2,3\}$. Fig. 4 shows $\mathcal{X}_{\parallel}$, $\mathcal{X}_{\perp}$, and their ratio; as $K\to 0$, the ratio tends to $1$, in agreement with the corollary and placing MRePU in the $K^{\star}=0$ universality class. With an appropriate choice of $C_{W}$, the kernel is kept from exploding or vanishing.
Figure 4: Parallel and perpendicular susceptibilities and their ratio for MRePU with $p=2$ (left) and $p=3$ (right), with $(C_{W},C_{b})=(1,0)$. The ratio tends to $1$ as $K\to 0$.
@@ 4.2 Experimental Results
Empirical Kernel and Four-point Vertex Behavior at Initialization As with the RePU case, we investigate the statistical properties of the neural network ensemble given by the initialization distribution. The experimental environment remains identical to that of the RePU Networks, with the only change being the activation function, which is now MRePU with $p=2$. The results for $C_{W}=1.0$ are shown in Fig. 5. As can be seen from the graph, the empirical kernel maintains a consistent scale across both random data and weight parameters. Additionally, we also examined the behavior of the Four-point vertex for MRePU. Eq. (5) can be simplified for a single input as follows:
@@
Fig. 6 illustrates how the empirical four-point vertex, calculated from an ensemble of 1000 models using Eq. (3), changes across layers (with input $x_{0}=(1,0)$). $\mathcal{X}_{\parallel}$ increases as the kernel grows, and since the kernel itself depends on $C_{W}$, the inductive change in the four-point vertex is determined by $C_{W}$ according to Eq. (10). As evidence of this, the left figure shows that the four-point vertex increases more steeply as $C_{W}$ becomes larger. We also investigated the behavior of the four-point vertex across different widths. The results indicate that with smaller widths, perturbative effects in Eq. (10) become more pronounced, leading to a more rapid increase in the four-point vertex, and overall, higher values are observed. The four-point vertex serves as an indicator of non-Gaussianity, which is expected to increase as more layers with non-linear activations are added. Moreover, since smaller widths imply stronger interactions between neurons, these experimental results align well with our theoretical predictions.
@@
Dynamics of Statistics During Training Now, we will examine how the statistics of the MRePU Network ensemble change under training dynamics. The experimental setup is the same as in the case of the RePU Network, and the number of hidden layers is set to 5. In this case, we constructed the training dataset in the same way as in the RePU activation case, but the number of samples is set to 1000. Although the theoretical approximation suggests setting $C_{W}=1.0$, meaningful learning was only achieved with a slightly lower value, possibly due to the perturbative effects in the MRePU Network. In our experiments, we set $C_{W}=0.8$. The dynamics of the kernel and four-point vertex are depicted in Fig. 7, while the comparison of the ensemble’s mean output against the target for the test dataset is shown in Fig. 8. Surprisingly, we observed that even with a vanilla FCN (Fully Connected Network) without considering architectures like ResNet, which are designed to prevent issues like the vanishing gradient, the deep neural network with 5 hidden layers still achieved meaningful learning using the MRePU function we designed. Additionally, as shown on the left side of Fig.  7, we also calculated the perpendicular susceptibility, which showed a tendency to approach 1 as training progressed. On the right side of the same figure, it can be seen that the ensemble’s four-point vertex decreases to 0 during initialization but tends to increase linearly with the layers as training progresses. An increasing four-point vertex indicates growing interactions between neurons, which implies that representation learning is occurring. This suggests that the deep vanilla MRePU network is indeed engaging in meaningful representation learning.
@@
Hyperparameter Criteria from Kernel Analysis Our kernel-based analysis yields concrete guidelines for configuring MRePU networks to achieve training stability. First, as implied by Corollary 9 and Eq. (4), MRePU belongs to the $K^{\star}=0$ universality class. Consequently, the input (and hence the induced kernel scale) must be sufficiently small for the kernel magnitude to remain steady as depth increases. Next, fixing $K$ at a suitably small value, Eq. (9) shows that choosing
$\begin{split}C_{W}\simeq\frac{1}{\sum_{n=2}^{2p+2}A_{n}(p)K^{\frac{n}{2}-1}I_{n}\Big(-\frac{1}{\sqrt{K}}\Big)}\simeq\frac{1}{\sum_{n=0}^{2p}B_{n}(p)K^{\frac{n}{2}}I_{n}\Big(-\frac{1}{\sqrt{K}}\Big)}.\end{split}$ (11)
guarantees a stable kernel at initialization in theory. We validate the practical effectiveness of these criteria on real-world datasets—MNIST and CIFAR-10—in Section 5.3.